What is the slope of a line passing through the points -4,-3 and 8,-9? Plz answer really need help.
Answer by Boreal(15235)

change in y/change in x is slope
go in the same direction for the y and x
so change in y is from -3 to -9, write it as -3-(-9)=-3+9=6
change in x is -4 to 8 or -4-8=-12
slope is 6/-12 or -1/2
you can go the other direction -9-(-3)=-6
and 8-(-4)=12
-6/12=-1/2
you have to put y over x and you MUST go in the same direction for y and x. It doesn't matter which direction, so long as BOTH are done that way.